OpenStudy (anonymous):

PLEASE EXPLAIN 2^3 - (2 x 3) + (-4) =

OpenStudy (anonymous):

SOLVE

OpenStudy (anonymous):

I FORGOT ALL THE RULES ECT. ON HOW TO DO IT

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Remember PEDMAS? 2^3-(2 *3)+(-4) so do everything in the parenthesis first (2*3) = 6 so you have 2^3-6-4 2^3 = 8 now you have 8-6-4 = -2 is your answer